# Cluster-Dags as Powerful Background Knowledge For Causal Discovery

### v1

@@ 3.2 Cluster-FCI
For brevity of notation, as in zhang2008completeness, we also introduce the "*" sympol to denote either an arrowhead, circle or tail. This will be used later for edge orientations, where some edge marks don’t matter for the orientation rule. For example if the edge $X*\hskip-5.0pt-\hskip-3.99994pt\circ Y$ is oriented as $X*\hskip-5.0pt\rightarrow Y$, the edge mark on the left stays whatever it was before the orientation. In the original definition of partial ancestral graphs (zhang2008completeness), undirected edges "-" or $\circ\hskip-1.79993pt-"$ are present in order to consider potential selection variables. In our work we assume non-existence of selection variables, so we omit these undirected edges here.
@@
27: Use rules R0-R4, R8-R10 of (zhang2008completeness) to orient as many edge marks as possible.
@@
Compatibility of MAGs (which are partial mixed graphs that satisfy ancestrality and maximality (zhang2008completeness)) with C-ADMGs follows directly from this definition, too. To summarize, a graph $G_{pm}^{\prime}$ being compatible with $G_{C}$ means it can be generated from an algorithm doing edge deletions and orientations on $G_{pm}$.
The graph resulting from Algorithm˜3 need not necessarily be ancestral yet, as the input C-ADMG, and thus also the output, may contain almost directed cycles. An almost directed cycle is defined as $X\leftrightarrow Y$ and $X\in an_{Y}^{G}$ (zhang2008completeness). A direct conversion of the C-ADMG to a MAG, e.g., following hu2020fasteralgorithmsmarkovequivalence, is undesirable, as the following example demonstrates. Consider the graph in Fig.˜2. The MAG $G_{M}$ over $\{X_{1},\ldots,X_{5}\}$ is ancestral, but the corresponding C-ADMG $G_{C}$ with $C=\{C_{1},C_{2},C_{3}\},C_{1}=\{X_{1},X_{2}\},C_{2}=\{X_{3}\},C_{3}=\{X_{4},X_{5}\}$ is not, due to the almost directed cycle $C_{3}\leftrightarrow C_{1}\rightarrow C_{2}\rightarrow C_{3}$. Even though there is an almost directed cycle in $G_{C}$, $G_{M}$ is compatible with $G_{C}$. If we were to transform this C-ADMG into a MAG, the edge $C_{1}\leftrightarrow C_{2}$ would change to $C_{1}\rightarrow C_{2}$. The edge between $X_{2},X_{5}$ would consequently inherit this orientation as $X_{2}\rightarrow X_{5}$ contradicting the correct edge type $X_{2}\leftrightarrow X_{5}$. Therefore, we only reorient almost directed cycles at the very end of C-FCI, ensuring it outputs a valid PAG without introducing false edge information.
@@
We now highlight some differences between Cluster-FCI and FCITiers. Cluster-FCI starts with the first cluster of the topological ordering and works its way down along the topological ordering, while FCITiers (andrews2020completeness, Alg. 1) starts with the last and works its way up. However, due to the nature of FCITiers, this direction does not matter as it is running a version of FCI on disjoint sets of edges (FCIExogenous). These disjoint edge sets are derived from the TBK. The resulting edges from FCIExogenous are then added to the overall graph. On the contrary, Cluster-FCI pre-processes the entire fully connected graph according to the C-ADMG to obtain a partial mixed graph. It then removes further edges along the topological ordering. Our proposed C-FCI Algorithm˜4 is sound, but not complete (proof in Section˜A.4).
@@ 5 Conclusion
Interesting future directions for future work include to apply C-DAG background knowledge to score-based methods (see Appendix˜D for some first thoughts), to combine C-DAGs with other types of prior knowledge such as pairwise background knowledge, or to include selection variables in the causal discover process as well. Since C-ADMGs can contain background knowledge that violates ancestrality required for FCI, it will also be interesting to investigate under which types of background knowledge an extended FCI version remains complete. Lastly, using summary causal graphs (ferreira2025identifyingC)—also allowing for cycles—instead of C-DAGs for improved causal discovery is also an interesting direction for follow up work.
@@ A.1 A C-DAG leads to an MPDAG
Definition 8 (CPDAG (completed partially directed graph), andersson1997characterization).
@@
Figure 5: Meek’s orientation rules (meek2013causalinferencecausalexplanation) (Figure from (fang2022representationcausalbackgroundknowledge)).
Definition 9 (Pairwise Causal Constraints, fang2022representationcausalbackgroundknowledge).
@@
Definition 10 (Restricted Markov equivalence class,fang2022representationcausalbackgroundknowledge).
@@
Definition 11 (Maximally partially directed acyclic graph (MPDAG), fang2022representationcausalbackgroundknowledge).
@@ A.2 Further definitions, theorems and additional explanations
Definition 14 (d-separation in C-DAGs,anand2023causal).
@@
Definition 15 (mns (minimal neighbor separator), gupta2023local).
@@
Proposition 1 (Restricting separating set via mns,gupta2023local).
@@
Definition 19 (Inducing path,zhang2008completeness).
@@
Definition 20 (Discriminating path,zhang2008completeness).
@@
Proposition 2 (Markov equivalence criterion for MAGs,zhang2008completeness).
@@
Definition 21 (Partial ancestral graph, zhang2008completeness).
@@
Definition 22 (Possible d-separating set, andrews2020completeness).
@@
Definition 23 (Orientation rules for FCI, zhang2008completeness).
@@ A.3 Soundness and completeness of Cluster-PC
Proof: Let $\hat{G}$ be the output of the C-PC algorithm and $\bar{G}$ the MPDAG of the restricted Markov equivalence class $G^{*},B]$, obtained by restricting the CPDAG output from PC with $B$. $G$ is the ground truth DAG. $V$ is the node set and $E_{\hat{G}},E_{\bar{G}},E_{G^{*}}$ their edge sets, respectively. First, we show that $\hat{G}$ has the same skeleton as $\bar{G}$. Let $X-Y\in E_{\bar{G}}$ be any edge in $\bar{G}$. This means $X,Y$ are not d-separable in $G$, and thus any compatible C-DAG $G_{C}$ will not put $X,Y$ into non-adjacent clusters. Also, no CI test performed in C-PC will remove this edge by the global Markov property ($\forall S:X\perp\!\!\!\perp_{G}Y|S\Rightarrow X\perp\!\!\!\perp Y|S$, thus $\forall S:X\not\!\perp\!\!\!\perp Y|S\rightarrow X\not\!\perp\!\!\!\perp_{G}Y|S$). On the other hand, let $X,Y$ be non-adjacent in $\bar{G}$, so $\exists S:X\perp\!\!\!\perp Y|S$. W.l.o.g. let $Y\notin de_{X}\cup nb_{X}^{+}$, then by Proposition˜1 $mns_{X}(Y)\subset pa_{X}$ and $X\perp\!\!\!\perp Y|mns_{X}(Y)$. Furthermore, $mns_{X}(Y)\subset pa_{X}\subset nch_{X}$ and for every $S^{\prime}\subset nch_{x}$, the CI test $X\stackrel{{\scriptstyle?}}{{\perp\!\!\!\perp}}Y|S^{\prime}$ is performed in C-PC. Thus the conditional independence $X\perp\!\!\!\perp Y|mns_{X}(Y)$ will be found and $X,Y$ are non-adjacent in $\hat{G}$, too. Second, we have to show that $\hat{G}$ and $\bar{G}$ have the same arrowheads. Due to the same skeleton, it is clear that they will also have the same unshielded colliders and same orientations due to Meek’s edge orientation rules. The only thing left to show that a) additional edge orientations coming from the C-DAG edges $E_{C}$ are the same in $\hat{G}$ and $\bar{G}$, as well as that b) edge orientations from using Meek’s orientation rules on the edges that partly come from a) are the same. Any directed edge $X\rightarrow Y\in E_{\bar{G}}$ coming from $B$ will also be directed in $\hat{G}$ due to Algorithm˜1. This then also leads to any directed edge $X\rightarrow Y\in E_{\bar{G}}$ coming from using orientation rules when combining CPDAG $G^{*}$ with $B$, also being oriented in the last steps of C-PC when Meek’s orientation rules are applied, so $X\rightarrow Y\in E_{\hat{G}}$. $\square$
@@ B.2 Parameters and additional tables from the simulation studies
DAG generation method Erdos-Renyi Erdos-Renyi, Erdos-Renyi Erdos-Renyi
@@ B.3 How compatible C-DAGs are generated
For Simulation 4 we use another method we call ’cdag-first’. This is because we wanted to enforce TBK on the cluster graphs. This method first generates an Erdos-Renyi graph for the clusters, for example of size three again. Then it generates nodes for each cluster so that they sum up to the desired node number, say ten again. Then the graph is built according to the generated cluster graph and nodes, and some edges from that graph are dropped out, that probability is influenced by the n_edges parameter. Since the C-DAG is generated first, we can exclude bidirected edges. And a fully connected version of this C-DAG does satisfy TBK.
@@ Appendix C Pairwise characterization of C-DAGs
Equivalence statement: $\Rightarrow"$: Let $G=(V,E)$ be a graph over the same variables $V$ with edges $E$ being compatible with C-DAG $G_{C}$. Furthermore let $bk(G_{C})$ be as defined in Eq.˜8. The task is to show that $bk(G_{C})$ is true for $G$, specifically, that any edge between any $X_{i},X_{j}$ satisfies $bk(G_{C})$. Take any $X_{i},X_{j}\in V$ and their respective clusters $X_{i}\in C_{i},X_{j}\in C_{j}$. $C_{i},C_{j}$ relate to each other in exactly one of the four ways described in (i)-(iv). Whatever the cluster relationship is on the left hand side of (i)-(iv), the proofs of (i)-(iv) above show that the edge between $X_{i},X_{j}$ satisfies the constraint on the right hand side of (i)-(iv). So the boolean pairwise combination $bk(G_{C})$ is true for $G$.
$\Leftarrow"$: Let $G_{C}$ be a C-DAG and $bk(G_{C})$ be true for DAG $G$. The task is to show that $G$ is compatible with $G_{C}$. $G=(V,E)$ is compatible with $G_{C}$, if none of its edges $E$ contradict the C-DAG edges $E_{C}$. Take any $X_{i},X_{j}$ and their corresponding clusters $X_{i}\in C_{i},X_{j}\in C_{j}$. The edge between $X_{i},X_{j}$ can take any of the three forms $X_{i}\rightarrow X_{j}$, $X_{i}\leftarrow X_{j}$ or $X_{i}\cancel{-}X_{j}$. W.l.o.g. (no need to consider the case $C_{i}\leftarrow C_{j}$ due to symmetry) the C-DAG restriction on the edge between $X_{i},X_{j}$ can take any of the forms in (i)-(iv).
@@ Appendix D Using C-DAGs for score-based and continuous optimization discovery algorithms
In this paper we studied constraint based causal discovery, but DAGs can also be estimated via a constrained optimization problem (chickering2002optimal; zheng2018dags). Such a problem typically admits a form like

### v2

@@ 3.2 Cluster-FCI
For brevity of notation, as in Zhang (2008b), we also introduce the “*” symbol to denote either an arrowhead, circle or tail. This will be used later for edge orientations, where some edge marks don’t matter for the orientation rule. For example if the edge $X*\hskip-5.0pt-\hskip-3.99994pt\circ Y$ is oriented as $X*\hskip-5.0pt\rightarrow Y$, the edge mark on the left stays whatever it was before the orientation. In the original definition of partial ancestral graphs (Zhang, 2008b), undirected edges “-” or $\circ\hskip-1.79993pt$-” are present in order to consider potential selection variables. In our work we assume non-existence of selection variables, so we omit these undirected edges here.
@@
27: Use rules R0-R4, R8-R10 of (Zhang, 2008b) to orient as many edge marks as possible.
@@
Compatibility of MAGs (which are partial mixed graphs that satisfy ancestrality and maximality (Zhang, 2008b)) with C-ADMGs follows directly from this definition, too. To summarize, a graph $G_{pm}^{\prime}$ being compatible with $G_{C}$ means it can be generated from an algorithm doing edge deletions and orientations on $G_{pm}$.
The graph resulting from Algorithm˜3 need not necessarily be ancestral yet, as the input C-ADMG, and thus also the output, may contain almost directed cycles. An almost directed cycle is defined as $X\leftrightarrow Y$ and $X\in an_{Y}^{G}$ (Zhang, 2008b). A direct conversion of the C-ADMG to a MAG, e.g., following Hu & Evans (2020), is undesirable, as the following example demonstrates. Consider the graph in Fig.˜2. The MAG $G_{M}$ over $\{X_{1},\ldots,X_{5}\}$ is ancestral, but the corresponding C-ADMG $G_{C}$ with $C=\{C_{1},C_{2},C_{3}\},C_{1}=\{X_{1},X_{2}\},C_{2}=\{X_{3}\},C_{3}=\{X_{4},X_{5}\}$ is not, due to the almost directed cycle $C_{3}\leftrightarrow C_{1}\rightarrow C_{2}\rightarrow C_{3}$. Even though there is an almost directed cycle in $G_{C}$, $G_{M}$ is compatible with $G_{C}$. If we were to transform this C-ADMG into a MAG, the edge $C_{1}\leftrightarrow C_{2}$ would change to $C_{1}\rightarrow C_{2}$. The edge between $X_{2},X_{5}$ would consequently inherit this orientation as $X_{2}\rightarrow X_{5}$ contradicting the correct edge type $X_{2}\leftrightarrow X_{5}$. Therefore, we only reorient almost directed cycles at the very end of C-FCI, ensuring it outputs a valid PAG without introducing false edge information.
@@
We now highlight some differences between Cluster-FCI and FCITiers. Cluster-FCI starts with the first cluster of the topological ordering and works its way down along the topological ordering, while FCITiers (Andrews et al., 2020, Alg. 1) starts with the last and works its way up. However, due to the nature of FCITiers, this direction does not matter as it is running a version of FCI on disjoint sets of edges (FCIExogenous). These disjoint edge sets are derived from the TBK. The resulting edges from FCIExogenous are then added to the overall graph. On the contrary, Cluster-FCI pre-processes the entire fully connected graph according to the C-ADMG to obtain a partial mixed graph. It then removes further edges along the topological ordering. Our proposed C-FCI Algorithm˜4 is sound, but not complete (proof in Section˜A.4).
@@ 5 Conclusion
Interesting future directions for future work include to apply C-DAG background knowledge to score-based methods (see Appendix˜D for some first thoughts), to combine C-DAGs with other types of prior knowledge such as pairwise background knowledge, or to include selection variables in the causal discover process as well. Since C-ADMGs can contain background knowledge that violates ancestrality required for FCI, it will also be interesting to investigate under which types of background knowledge an extended FCI version remains complete. Lastly, using summary causal graphs (Ferreira & Assaad, 2025b)—also allowing for cycles—instead of C-DAGs for improved causal discovery is also an interesting direction for follow up work.
Acknowledgments
This work has been supported by the German Federal Ministry of Education and Research (Grant: 01IS24082).
@@ A.1 A C-DAG leads to an MPDAG
Definition 8 (CPDAG (completed partially directed graph), Andersson et al., 1997).
@@
Figure 5: Meek’s orientation rules (Meek, 1995) (Figure from (Fang et al., 2025)).
Definition 9 (Pairwise Causal Constraints, Fang et al., 2025).
@@
Definition 10 (Restricted Markov equivalence class,Fang et al., 2025).
@@
Definition 11 (Maximally partially directed acyclic graph (MPDAG), Fang et al., 2025).
@@ A.2 Further definitions, theorems and additional explanations
Definition 14 (d-separation in C-DAGs,Anand et al., 2023).
@@
Definition 15 (mns (minimal neighbor separator), Gupta et al., 2023).
@@
Proposition 1 (Restricting separating set via mns,Gupta et al., 2023).
@@
Definition 19 (Inducing path,Zhang, 2008b).
@@
Definition 20 (Discriminating path,Zhang, 2008b).
@@
Proposition 2 (Markov equivalence criterion for MAGs,Zhang, 2008b).
@@
Definition 21 (Partial ancestral graph, Zhang, 2008b).
@@
Definition 22 (Possible d-separating set, Andrews et al., 2020).
@@
Definition 23 (Orientation rules for FCI, Zhang, 2008b).
@@ A.3 Soundness and completeness of Cluster-PC
Proof: Let $\hat{G}$ be the output of the C-PC algorithm and $\bar{G}$ the MPDAG of the restricted Markov equivalence class $G^{*},B]$, obtained by restricting the CPDAG output from PC with $B$. $G$ is the ground truth DAG. $V$ is the node set and $E_{\hat{G}},E_{\bar{G}},E_{G^{*}}$ their edge sets, respectively. First, we show that $\hat{G}$ has the same skeleton as $\bar{G}$. Let $X-Y\in E_{\bar{G}}$ be any edge in $\bar{G}$. This means $X,Y$ are not d-separable in $G$, and thus any compatible C-DAG $G_{C}$ will not put $X,Y$ into non-adjacent clusters. Also, no CI test performed in C-PC will remove this edge by the global Markov property ($\forall S:X\perp\!\!\!\perp_{G}Y|S\Rightarrow X\perp\!\!\!\perp Y|S$, thus $\forall S:X\not\!\perp\!\!\!\perp Y|S\rightarrow X\not\!\perp\!\!\!\perp_{G}Y|S$). On the other hand, let $X,Y$ be non-adjacent in $\bar{G}$, so $\exists S:X\perp\!\!\!\perp Y|S$. Without loss of generality, let $Y\notin de_{X}\cup nb_{X}^{+}$, then by Proposition˜1 $mns_{X}(Y)\subset pa_{X}$ and $X\perp\!\!\!\perp Y|mns_{X}(Y)$. Furthermore, $mns_{X}(Y)\subset pa_{X}\subset nch_{X}$ and for every $S^{\prime}\subset nch_{x}$, the CI test $X\stackrel{{\scriptstyle?}}{{\perp\!\!\!\perp}}Y|S^{\prime}$ is performed in C-PC. Thus the conditional independence $X\perp\!\!\!\perp Y|mns_{X}(Y)$ will be found and $X,Y$ are non-adjacent in $\hat{G}$, too. Second, we have to show that $\hat{G}$ and $\bar{G}$ have the same arrowheads. Due to the same skeleton, it is clear that they will also have the same unshielded colliders and same orientations due to Meek’s edge orientation rules. The only thing left to show that a) additional edge orientations coming from the C-DAG edges $E_{C}$ are the same in $\hat{G}$ and $\bar{G}$, as well as that b) edge orientations from using Meek’s orientation rules on the edges that partly come from a) are the same. Any directed edge $X\rightarrow Y\in E_{\bar{G}}$ coming from $B$ will also be directed in $\hat{G}$ due to Algorithm˜1. This then also leads to any directed edge $X\rightarrow Y\in E_{\bar{G}}$ coming from using orientation rules when combining CPDAG $G^{*}$ with $B$, also being oriented in the last steps of C-PC when Meek’s orientation rules are applied, so $X\rightarrow Y\in E_{\hat{G}}$. $\square$
@@ B.2 Parameters and additional tables from the simulation studies
DAG generation method Erdős–Rényi Erdős–Rényi Erdős–Rényi Erdős–Rényi
@@ B.3 How compatible C-DAGs are generated
For Simulation 4 we use another method we call ’cdag-first’. This is because we wanted to enforce TBK on the cluster graphs. This method first generates an Erdős–Rényi graph for the clusters, for example of size three again. Then it generates nodes for each cluster so that they sum up to the desired node number, say ten again. Then the graph is built according to the generated cluster graph and nodes, and some edges from that graph are dropped out, that probability is influenced by the n_edges parameter. Since the C-DAG is generated first, we can exclude bidirected edges. And a fully connected version of this C-DAG does satisfy TBK.
@@ Appendix C Pairwise characterization of C-DAGs
Equivalence statement: $\Rightarrow$”: Let $G=(V,E)$ be a graph over the same variables $V$ with edges $E$ being compatible with C-DAG $G_{C}$. Furthermore let $bk(G_{C})$ be as defined in Eq.˜8. The task is to show that $bk(G_{C})$ is true for $G$, specifically, that any edge between any $X_{i},X_{j}$ satisfies $bk(G_{C})$. Take any $X_{i},X_{j}\in V$ and their respective clusters $X_{i}\in C_{i},X_{j}\in C_{j}$. $C_{i},C_{j}$ relate to each other in exactly one of the four ways described in (i)-(iv). Whatever the cluster relationship is on the left hand side of (i)-(iv), the proofs of (i)-(iv) above show that the edge between $X_{i},X_{j}$ satisfies the constraint on the right hand side of (i)-(iv). So the boolean pairwise combination $bk(G_{C})$ is true for $G$.
$\Leftarrow$”: Let $G_{C}$ be a C-DAG and $bk(G_{C})$ be true for DAG $G$. The task is to show that $G$ is compatible with $G_{C}$. $G=(V,E)$ is compatible with $G_{C}$, if none of its edges $E$ contradict the C-DAG edges $E_{C}$. Take any $X_{i},X_{j}$ and their corresponding clusters $X_{i}\in C_{i},X_{j}\in C_{j}$. The edge between $X_{i},X_{j}$ can take any of the three forms $X_{i}\rightarrow X_{j}$, $X_{i}\leftarrow X_{j}$ or $X_{i}\cancel{-}X_{j}$. Without loss of generality (no need to consider the case $C_{i}\leftarrow C_{j}$ due to symmetry), the C-DAG restriction on the edge between $X_{i},X_{j}$ can take any of the forms in (i)-(iv).
@@ Appendix D Using C-DAGs for score-based and continuous optimization discovery algorithms
In this paper we studied constraint based causal discovery, but DAGs can also be estimated via a constrained optimization problem (Chickering, 2002; Zheng et al., 2018). Such a problem typically admits a form like